# Well-posedness of a generalized Stokes operator on domains with cylindrical ends via layer-potentials

### 11 May 2026 v1

@@ 11.2. The non-homogeneous Stokes system
A useful consequence of Theorem 11.4 is the following corollary.
@@
Let the assumptions of Theorem 1.1 hold. Then, for every $m\in\mathbb{Z}_{+}$, there exists a constant ${\mathcal{C}}_{m}\in(0,\infty)$ such that, for any $\boldsymbol{f}\in H^{m+1/2}(\Gamma;TM)$ and any $\boldsymbol{h}\in H^{m-1}(\Omega;TM)$, the non-homogeneous Dirichlet problem (11.2) has a unique solution $U=(\boldsymbol{u}\ \ p)^{\top}\in H^{m+1}(\Omega;TM)\oplus H^{m}(\Omega)$ and this solution satisfies
@@
Let assumptions of Theorem 1.1 hold. Let $\mathcal{N}_{\rm{ST}}:H^{3/2}(\partial\Omega;TM)\to H^{1/2}(\partial\Omega;TM)$ be the Dirichlet-to-Neumann operator defined as follows (see [131, p.37] in the case of the Laplace operator on a compact manifold). For $\boldsymbol{f}\in H^{3/2}(\partial\Omega;TM)$ arbitrary, let $U:=(\boldsymbol{u}\ \ p)^{\top}\in H^{2}(\Omega;TM)\oplus H^{1}(\Omega)$ be the unique solution of the Dirichlet problem (11.1) (that is, $\boldsymbol{\Xi}_{V,V_{0}}U\,=\,0$ in $\Omega$ and $\boldsymbol{u}|_{\Gamma}\,=\,\boldsymbol{f}$ on $\Gamma:=\partial\Omega$), see Theorem 1.1. Then we define
@@
The next result shows that there is no jump of the conormal derivative $\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{D}_{\rm{ST}}(\boldsymbol{f})$ across $\Gamma:=\partial\Omega$ (see [131, Proposition 11.4] in the case of the Laplace operator).
@@ 12.1. The Dirichlet problem for the generalized Navier-Stokes system
A key role in the analysis of the nonlinear Dirichlet problem for the Navier-Stokes system on a domain with cylindrical ends is played by estimates of the product of two functions in suitable Sobolev spaces. An example is the following result, which is well known in $\mathbb{R}^{n}$ (see e.g. [17, Theorem 7.3, Theorem 8.2], [114, §4.5.2, Theorem], and [90, Lemma 8.1]).
@@
We notice that the same argument works for manifolds with bounded geometry. Also, notice that the result of Lemma 12.1 extends to Sobolev spaces on sections of vector bundles over $M$ (see e.g. [114, §4.2.2, Proposition] in the Euclidean setting), and the next lemma is a consequence of this result (see [39, Lemma 7.5] in the case of a Lipschitz domain on a compact manifold).
@@
(see [114, §4.2.2, Proposition] in $\mathbb{R}^{n}$). This result extends also to vector bundles and, thus, there exists a constant $C>0$ (depending only on $n$, $m$ and the bounded geometry of $M$) such that the estimate (12.1) holds for all $\boldsymbol{u},\boldsymbol{v}\in H^{m+1}(\Omega;TM)$.
@@
Let us now prove the well-posedness of the Dirichlet problem for the Navier-Stokes system under the assumption of small data (see Theorem 7.3 in [39] in the case of a $C^{1}$, or even Lipschitz domain on a compact Riemannian manifold).
@@ 12.2. The case $M$ closed
    There are no restrictions “or conditions” at infinity on our operators, functions, and vector bundles. For instance, $\mathcal{C}^{\infty}_{\operatorname{inv}}$ is replaced with $\mathcal{C}^{\infty}$, all open subsets and all vector bundles are compatible, and $\Psi_{\operatorname{ess}}^{m}$ is replaced with $\Psi^{m}$.
@@
    The restuls about the limit at infinity $\mathcal{R}_{\infty}$} do not make sense anymore (and are not needed). They have to be ignored in the closed case.
@@
The case $M$ closed was considered in great detail in [69], so we will not go through the exercice of performing all the modifications abour. Instead, we recall here only two results from that paper. The first result describes the kernel of $\boldsymbol{\Xi}$ and does not follow from the results of this paper (directly):
@@
Let us assume that $M$ is a closed manifold, that $V,V_{0}$ are smooth and non-negative and that $M$ is a smooth, compact manifold without boundary (i.e., a closed manifold). Then $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}:H^{1}(M;TM)\oplus L^{2}(M)\to H^{-1}(M;TM)\oplus L^{2}(M)$ is a self-adjoint Fredholm operator. Let $\mathcal{N}\subset\mathcal{C}^{\infty}(M;TM\oplus\mathbb{C})$ be defined by
@@
Then the kernel of $\boldsymbol{\Xi}_{V,V_{0}}$ is given by $\ker\boldsymbol{\Xi}_{V,V_{0}}=\mathcal{N}$. Moreover, $\boldsymbol{\Xi}$ has a (unique) Moore-Penrose pseudoinverse $\boldsymbol{\Xi}^{(-1)}\in\Psi_{\operatorname{cl}}^{-\boldsymbol{s}-\boldsymbol{t}}(M;TM\oplus\mathbb{C})$, with $\boldsymbol{s}=\boldsymbol{t}=(1,0)$, whose image is $\mathcal{N}^{\perp}$.
@@
Any two solutions $U_{1}$ and $U_{2}$ of this Dirichlet problem differ by a constant scalar field: $U_{2}-U_{1}=(0\ \ \ c)^{\top}$, $c\in\mathbb{C}$. Also, any of the following two formulas provides a solution of this Dirichlet problem:
@@
Moreover, there exists a constant $C_{m}\geq 0$, independent of $\boldsymbol{h}$, such that all solutions $U=(\boldsymbol{u}\ \ p)^{\top}$ satisfy

### 28 May 2026 v2

@@ 11.2. The non-homogeneous Stokes system
Let us now consider the case $r=0$ in Equation (11.2). Then a consequence of Theorem 11.4 that will be used in what follows is the following corollary.
@@
Let us further assume in Theorem 11.4 that $r=0$. Then, for every $m\in\mathbb{Z}_{+}$, there exists a constant ${\mathcal{C}}_{m}\in(0,\infty)$ such that, for any $\boldsymbol{f}\in H^{m+1/2}(\Gamma;TM)$ and any $\boldsymbol{h}\in H^{m-1}(\Omega;TM)$, the non-homogeneous Dirichlet problem (11.2) has a unique solution $U=(\boldsymbol{u}\ \ p)^{\top}\in H^{m+1}(\Omega;TM)\oplus H^{m}(\Omega)$ and this solution satisfies
@@
Let assumptions of Theorem 1.1 hold. Let $\mathcal{N}_{\rm{ST}}:H^{3/2}(\partial\Omega;TM)\to H^{1/2}(\partial\Omega;TM)$ be the Dirichlet-to-Neumann operator defined as follows (see [155, p.37] in the case of the Laplace operator on a compact manifold). For $\boldsymbol{f}\in H^{3/2}(\partial\Omega;TM)$ arbitrary, let $U:=(\boldsymbol{u}\ \ p)^{\top}\in H^{2}(\Omega;TM)\oplus H^{1}(\Omega)$ be the unique solution of the Dirichlet problem (11.1) (that is, $\boldsymbol{\Xi}_{V,V_{0}}U\,=\,0$ in $\Omega$ and $\boldsymbol{u}|_{\Gamma}\,=\,\boldsymbol{f}$ on $\Gamma:=\partial\Omega$), see Theorem 1.1. Then we define
@@
The next result shows that there is no jump of the conormal derivative $\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{D}_{\rm{ST}}(\boldsymbol{f})$ across $\Gamma:=\partial\Omega$ (see [155, Proposition 11.4] in the case of the Laplace operator).
@@ 12.1. The Dirichlet problem for the generalized Navier-Stokes system
A key role in the analysis of the nonlinear Dirichlet problem for the Navier-Stokes system on a domain with cylindrical ends is played by estimates of the product of two functions in suitable Sobolev spaces. An example is the following result, which is well known in $\mathbb{R}^{n}$ (see e.g. [21, Theorem 7.3, Theorem 8.2], [137, §4.5.2, Theorem], and [111, Lemma 8.1]).
@@
We notice that the same argument works for manifolds with bounded geometry. Also, notice that the result of Lemma 12.1 extends to Sobolev spaces on sections of vector bundles over $M$ (see e.g. [137, §4.2.2, Proposition] in the Euclidean setting), and the next lemma is a consequence of this result (see [49, Lemma 7.5] in the case of a Lipschitz domain on a compact manifold).
@@
(see [137, §4.2.2, Proposition] in $\mathbb{R}^{n}$). This result extends also to vector bundles and, thus, there exists a constant $C>0$ (depending only on $n$, $m$ and the bounded geometry of $M$) such that the estimate (12.1) holds for all $\boldsymbol{u},\boldsymbol{v}\in H^{m+1}(\Omega;TM)$.
@@
Let us now prove the well-posedness of the Dirichlet problem for the Navier-Stokes system under the assumption of small data (see also [17, 49, 83, 111] for results in bounded Lipschitz domains).
@@ 12.2. The case $M$ closed
    There are no restrictions “or conditions” at infinity on our operators, functions, and vector bundles. For instance, $\mathcal{C}^{\infty}_{\operatorname{inv}}$ is replaced with $\mathcal{C}^{\infty}$, all open subsets and all vector bundles are compatible, and $\Psi_{\operatorname{inv}}^{m}$ and $\Psi_{\operatorname{ess}}^{m}$ are replaced with $\Psi^{m}$, $m\in\mathbb{Z}\cup\{\pm\infty\}$.
@@
    The results about the limit at infinity $\mathcal{R}_{\infty}$} do not make sense anymore (and are not needed). They have to be ignored in the closed case.
@@
The case $M$ closed was considered in great detail in [89], so we will not go through the exercice of performing all the modifications needed here. Instead, we recall here only two results from that paper that do not follow directly from the results of this paper. The first result describes the kernel of $\boldsymbol{\Xi}$. The second result formulates the well-posendess of the Stokes operator.
@@
Let us assume that $V,V_{0}$ are smooth and non-negative and that $M$ is a smooth, compact manifold without boundary (i.e., a closed manifold). Then $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}:H^{1}(M;TM)\oplus L^{2}(M)\to H^{-1}(M;TM)\oplus L^{2}(M)$ is a self-adjoint Fredholm operator. Let $\mathcal{N}\subset\mathcal{C}^{\infty}(M;TM\oplus\mathbb{C})$ be defined by
@@
Then $\mathcal{N}$ is the kernel of $\boldsymbol{\Xi}_{V,V_{0}}$ (i.e., $\ker\boldsymbol{\Xi}_{V,V_{0}}=\mathcal{N}$). Moreover, $\boldsymbol{\Xi}$ has a (unique) Moore-Penrose pseudoinverse $\boldsymbol{\Xi}^{(-1)}\in\Psi_{\operatorname{cl}}^{-\boldsymbol{s}-\boldsymbol{t}}(M;TM\oplus\mathbb{C})$, with $\boldsymbol{s}=\boldsymbol{t}=(1,0)$, whose image is $\mathcal{N}^{\perp}$.
@@
Any two solutions $U_{1}$ and $U_{2}$ of this Dirichlet problem differ by a constant scalar field: $U_{2}-U_{1}=(0\ \ \ c)^{\top}$, $c\in\mathbb{C}$. Moreover, there exists a constant $C_{m}\geq 0$, independent of $\boldsymbol{h}$, such that all solutions $U=(\boldsymbol{u}\ \ p)^{\top}$ satisfy
Moreover, using layer potentials, we obtain explicit formulas for the solution. Indeed, any of the following two formulas provides a solution of the Dirichlet problem for $\boldsymbol{\Xi}_{V,0}$ in the above theorem:
@@
If, however, $V_{0}\not\equiv 0$ on $\Omega$ in the above theorem, we obtain the uniqueness of the solution $U$, as in this paper.